I made some efforts to fix the situation of SRM but it is still bad.
The POS motion wasn't well damped. Something is wrong either (maybe both) sensing part or actuation part.
I am going to check the sensing matrix with the new free swinging spectra (#5690)
When I was trying to lock SRMI I found that the fringes observed at the AS camera didn't show spatial higher order modes, which is good.
So I thought the SRM suspension became quiet, but it actually wasn't. Because the RMS monitor of the SRM OSEMs already went to about 30 counts.
At the same time the opelev error signals were well suppressed. It means some DOFs which were insensitive to the oplev were ringing up, namely POS and SIDE.
According to the LSC error signal and the ASDC signal, I believe that the POS was going wild (although I didn't check the OSEM spectra).
+ Readjusted the f2a filters (see the attachment).
+ Tried to eliminate a coupling between the POS and SIDE drives by tweaking the output matrix.
=> In order to eliminate the coupling from the POS drive to SIDE sensor, I had to put a comparable factor into an element.
So it might be possible that the POS sensor was showing the SIDE signal and vice versa.
In order to check it I left SRM free swinging (#5690).
The main reason why I couldn't lock DRMI was that the suspensions were touchy and especially the SRM suspension wasn't good.